OpenStudy (anonymous):

What is the momentum of an object that has moved 5 meters in a period of 16 seconds if that object has a mass of 32 kg?

OpenStudy (xkehaulanix):

I believe this is a physics question? But okay... The momentum of an object is equal to its mass times velocity, or momentum = mass * velocity The velocity in turn is equal to distance divided by time, so velocity = 5/16 m/s So multiply the velocity (5/16) by the mass (32 kg) 5/16 x 32 = 10 I believe that's right? I haven't taken physics in a while. Meters per second and kilograms are the proper units, aren't they?

OpenStudy (anonymous):

ok ill take your word for it thanks a bunch!

OpenStudy (anonymous):

hey kehaul do you think you could help me with one more?: A 30 kg sled moving at 15m/s accels down a hill to reach a speed of 20m/s. It takes 4.3 secs to reach the bottom of the hill. What is the sled's change in momentum?

OpenStudy (xkehaulanix):

Er....I'm really not sure about this one, but the change in momentum should be equal to the difference between the final and initial momentum. So using the formula for momentum, you should get this: Final_momentum = 30 kg * 20 m/s = 600 Initial_momentum = 30 kg * 15 m/s = 450 Change_in_momentum = 150 kg*m/s Again, I'm not really sure. Physics is a little iffy for me.

OpenStudy (anonymous):

ya thats what i got but i wasnt a 100% sure. thanks again!!